OpenStudy (anonymous):

Solve the differential equation dq/dz = 2 + sin(z) with initial condition q(π) = 2π

OpenStudy (astrophysics):

Dibs

OpenStudy (anonymous):

you got it!

OpenStudy (astrophysics):

\[dq = 2+\sin(z) dz \implies \int\limits dq = \int\limits (2+\sin(z)) dz\]

OpenStudy (astrophysics):

separable*

OpenStudy (astrophysics):

Integrate it now

OpenStudy (anonymous):

alright so i got the portion with 2z-cos(z)

OpenStudy (anonymous):

However im missing something

OpenStudy (astrophysics):

\[q = 2z-cosz+C\]

OpenStudy (astrophysics):

Solve for the constant C using your initial conditions

OpenStudy (anonymous):

Thanks!

OpenStudy (astrophysics):

Yw :)

OpenStudy (anonymous):

Im probably going to have a lot more questions coming too I hate finals!

OpenStudy (astrophysics):

Haha, awesome! :P

OpenStudy (astrophysics):

separable are one of my fav questions